OpenStudy (anonymous):

Find the equation of the ellipse given the foci (plus or minus 4, 0) ; minor axis is 6

OpenStudy (anonymous):

@nincompoop @Directrix @mathmate

OpenStudy (mathmate):



OpenStudy (mathmate):

The definition of the ellipse: a=major axis, b=minor axis, c=semi-focal distance. By Pythagoras theorem, \(a^2=b^2+c^2.\) The corresponding equation of the ellipse is \(\Large \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\)

OpenStudy (anonymous):

so c^2=a^2-b^2 16=a^2-36 a^2=16+36 a^2=52 then the equation is \[\frac{ x^2 }{ 52 }+\frac{ y^2 }{ 36 }=1\]

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

hi!!

OpenStudy (freckles):

looks good sqrt(36)<sqrt(52) so the minor axis is sqrt(36) and the foci will lay on the major axis and so we are basically given the major axis is the x-axis because (-4,0) and (4,0) are points that lay on the x-axis ...